OpenStudy (anonymous):

Need help medal for you Factor the expression, show the work 8x4-36x3 4x3((2x)+4x3(-9) 4x3(2x-9)

OpenStudy (anonymous):

is this breakdown correct

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

did i miss anything ?

OpenStudy (skullpatrol):

$$ 8x^4-36x^3$$ $$=4x^3((2x)+4x^3(-9)$$ $$=4x^3(2x-9)$$

OpenStudy (skullpatrol):

you missed: equals signs and ^ for exponents

OpenStudy (anonymous):

ok ty im trying hard to figure out the correct way to break it down. can u check another for me

OpenStudy (anonymous):

28x2-23x-15 = 28x2-35+12x-15 = 7x (4x-5) 3(4x-5)) =(7x+3)(4x-5)